OpenStudy (tayaiv):

Expand this: √(x-1)² + (y-2)²

OpenStudy (tayaiv):

Apparently, the next step is: x² + y² = x² - 2x + 1 + y² -4y + 4 I'm wondering where the -2x and the -4y came from.

OpenStudy (tayaiv):

@Directrix @DanJS Could you please help if you can?

OpenStudy (anonymous):

(x-1)^2 = (x-1)(x-1) = x^2 - x - x + 1 = x^2 -2x +1 (y-2)^2 = (y-2)(y-2) = y^2 - 2y - 2y +4 = y^2 - 4y + 4 General formula: (a - b)^2 = a^2 - 2ab + b^2